OpenStudy (anonymous):

How do I write (5,8) and (0,-5) in standard form?

Directrix (directrix):

First, write the equation of the line determined by the two points (5,8) and (0,-5). Get the slope which is the change in y over change in x or (y1 - y2)/(x1 - x2)/ In this case, that value is (8 - (-5)) / (5 - 0) = 13/5 which is designated by m. Using the point-slope form of a line: y - y1 = m(x - x1), the equation of the line becomes y - (-5) = (13/5) (x - 0) where m = 13/5 and (x1, y1) = (0.-5). To continue, y + 5 = (13/5) x - (13/5)(0) = y + 5 = (13/5) x. y + 5 = (13/5) x must be written in standard form. So, next comes ....

Directrix (directrix):

Standard Form: the standard form of a line is in the form Ax + By = C where A is a positive integer, and B, and C are integers. y + 5 = (13/5) x Multiply each term by 5. 5y + 25 = 13x 13x - 5y = 25 13x + (-5)y = 25 where A = 13, B = -5, and C = 25 -->* Check my work, please.